Use the substitution method to solve the system of equations.
2x - 5y = -14
x - 2y = -5

Answer by rfer(16322)

2x-5y=-14
x-2y=-5 x=2y-5
----------------
2(2y-5)-5y=-14
4y-10-5y=-14
-y=-14+10
-y=-4
y=+4
----------------
2x-5*4=-14
2x-20=-14
2x=-14+20
2x=6
x=6/2
x=3
-----------
x-2y=-5
3-2*4=-5
3-8=-5